If a vehicle travels 1320 feet in seven seconds from a dead stop , how long does it take for that same vehicle to travel 1313 feet from a dead stop ? Thank you
Found 2 solutions by Alan3354, richard1234:
Answer by Alan3354(69443)

If a vehicle travels 1320 feet in seven seconds from a dead stop , how long does it take for that same vehicle to travel 1313 feet from a dead stop ?
--------------
You have to assume that the acceleration is uniform, or, if not, give detail on it.
-----
s = at^2/2 (s = distance, a = acceleration)
1320 = a*49/2
a = 2640/49 ft/sec/sec
-------------
1313 = (2640/49)*t^2/2
t^2 = 2626*49/2640
t =~ 6.981415 seconds

Answer by richard1234(7193)

The best answer is "slightly under 7 seconds" because the vehicle's speed is almost certainly not constant, and you have to factor acceleration into the solution. Assuming constant acceleration a, you may use to solve for the acceleration, then plug back in.